# Truck Parking Usage Prediction with Decomposed Graph Neural Networks

### v1

@@ IV-F Discussion
The efficiency of our models in terms of their ability to achieve superior performance stems from a reduced number of overlaps among reference nodes. Specifically, in a connected graph with self-loops, each node is calculated $n*l$ times, where $n$ expresses the number of nodes and $l$ represents the number of neighboring nodes. However, these overlaps hinder the learning of the graph structure since a node requires more computation to refer to more nodes. In contrast, our model addresses this issue by creating subgraphs and reducing the overlap to $(v/r)*l$ for each spatial module, where $r$ denotes the number of subgraphs. As a result, the model can now prioritize the crucial truck parking feature in the region through simple calculations.
@@
From the experiment results, it is evident that the Regional Decomposition technique outperforms both the connected and randomly connected graph configurations. Also, when compared to CST-GCN, the RegT-GCN demonstrates a significant performance advantage. This suggests that the concatenation of hidden vectors plays a crucial role in sharing subgraph features, highlighting the importance of employing graph neural networks in conjunction with pre-clustering graphs for superior performance. As a potential future research, we propose exploring the concept of decomposing by highway functions, which relates closer to truck drivers’ behaviors.
@@
According to the analysis of previous research [12], truck drivers tend to use consecutive parking spaces when one of them is fully occupied. Consequently, considering cooperative node interactions within regions yields better outcomes. Hence, the competitive performance of the Regional Decomposition approach in comparison to a single connected graph indicates that our framework can achieve strong performance simply by decomposing the input graph.
In our future research, we will investigate the limits of information preservation on our models by incorporating deeper layers and exploring the potential for extending prediction time horizons to their maximum allowable limits. Furthermore, we acknowledge that the relatively inferior results observed in longer time horizons may stem from limitations in computation power for handling larger training datasets. Therefore, we anticipate that utilizing larger training datasets will yield performance improvements for our models.
@@
Lastly, it is necessary to design an end-to-end framework capable of automatically decomposing graphs and integrating spatio-temporal modules for inductive scenarios. This will enable us to apply the models to previously unseen graphs, such as those outside the scope of participating states in MAASTO while maintaining consistent and robust performance. These future directions hold the potential to further advance the capabilities and applicability of our proposed framework.
@@ V Conclusion
In this paper, we propose the Regional Temporal Graph Neural Network (RegT-GCN), a state-wide truck parking usage prediction framework, leveraging topological structures of truck parking site distributions. The RegT-GCN comprises a Regional Decomposition approach to learn the geographical characteristics efficiently and a spatio-temporal module to capture the temporal dependencies on graphs. Experiments using the field data show that the proposed models outperform other baseline models. Decomposed graphs into regional oriented subgraphs result in better prediction capability and consistency of inferences. The study also encompasses a demonstration of predicting usages through the empirical observation of non-informative loss, facilitated by the integration of the Gated Recurrent Unit module, which adeptly adapts to value fluctuations. Through extensive experiments focused on graph connectivity, this paper validates that incorporating regional relationships substantially enhances the comprehension of the structural dynamics within truck parking site location graphs.

### v2

@@ IV-G Discussion
The superior performance of our models stems from a reduced number of overlaps among reference nodes. Specifically, in a connected graph with self-loops, each node is calculated $N*l$ times, where $N$ expresses the number of nodes and $l$ represents the number of neighbors being sampled for each node. However, these overlaps hinder the learning of the graph structure since a node requires more computation to refer to more nodes. In contrast, our model addresses this issue by creating subgraphs and reducing the overlap to $N^{r}*l^{r}$ for each spatial module, where $N^{r}$ and $l^{r}$ denote the number of nodes and neighboring nodes in subgraphs. As a result, the model can prioritize the crucial truck parking feature in the region through simple calculations.
@@
The experiment results show that both the regional and random decomposition technique outperform the non-decomposed method, T-GCN. This suggests that the concatenation of hidden vectors plays a crucial role in sharing subgraph features, highlighting the importance of employing graph neural networks in conjunction with pre-clustering graphs for superior performance.
@@
At the same time, the connectivity of the decomposed graph needs to be taken into account. According to the analysis of previous research [12], truck drivers tend to use consecutive parking sites when one of them is fully occupied. Consequently, considering cooperative node interactions within regions yields better outcomes. Hence, the competitive performance of the Regional Decomposition approach in comparison to a single or random connected graph indicates that our framework can achieve strong performance simply by decomposing the input graph.
As a potential future research, we propose exploring the concept of decomposing by highway functions, which is more closely related to the behaviors of truck drivers. Graph connectivity also needs to be reconfigured aligned with truck drivers decision making process, which they tend to seek truck parking sites on the same or close highway.
@@
The investigation of the limits of over-smoothing [33] on our models is also important. By incorporating deeper layers and exploring the potential for extending prediction time horizon, we will find our model’s maximum allowable performance limits. Lastly, it is necessary to design an end-to-end framework capable of automatically decomposing graphs and integrating spatio-temporal modules for inductive scenarios. This will enable us to apply the models to previously unseen graphs, such as those outside the scope of participating states in MAASTO while maintaining consistent and robust performance. These future directions hold the potential to further advance the capabilities and applicability of our proposed framework.
@@ V Conclusion
In this paper, we propose the Regional Temporal Graph Neural Network (RegT-GCN), a state-wide truck parking usage prediction framework, leveraging topological structures of truck parking site distributions. The RegT-GCN comprises a Regional Decomposition approach to learn the geographical characteristics efficiently and a spatio-temporal module to capture the temporal dependencies on graphs. Experiments using our original field dataset show that the proposed models outperform other baseline models. Decomposed graphs into regional-oriented subgraphs result in better prediction capability and consistency of inferences. The study also encompasses a demonstration of predicting usages through the empirical observation of non-informative loss, facilitated by the integration of the Gated Recurrent Unit module, which adeptly adapts to value fluctuations. Furthermore, with extensive experiments focused on graph connectivity, this paper validates that incorporating regional relationships of truck parking enhances the comprehension of the structural dynamics of truck parking behavior. In future work, we will explore better graph connectivity based on highway functions in an end-to-end manner.